1/2

Ирина Саверьевна Алексеева ,  Ирина Саверьевна Алексеева ,  Ирина Саверьевна Алексеева ,
Oда Самарканду  Ode a Samarcanda  Ode to Samarkand
tradukita de Carlo Minnaja  tradukita de Malcolm Jones

Пою тебя, о, Самарканд! Твой чудный облик восхваляю. И твоё сердце - Регистан - Я вечно юным называю.

Argiope-informo:

La aspekto de tiu ĉi versio estas barita.

Supozeble la aŭtoro decidis tion.

I sing of you, oh, Samarkand’!

Your wondrous features win my praises. And your dear heart - the Registan - seems young to me and never ages.

Его биенье Улугбек Предчувствовал под небом звёздным, О, Самарканд! Твой долгий век Величием Тимура создан!

Its steady beating Ulughbek

could sense beneath the constellations. Oh, Samarkand! Immortal pick of mighty Tamerlan’s creations.

Твои святые медресе - Приют искусств и просвещенья, Они стоят во всей красе Столетья превратив в мгновенья.

Madrasses are your sacred schools, which cultivate the arts and learning. Their beauty overcomes time’s rules by centuries into moments turning.

Здесь Алишера Навои Поэзию - любовь и силу - Впитали жители твои Она их счастьем вдохновила.

Here the great Navoi’s verse, so full of love and strong emotion, is cherished by your populace, made happy by its inspiration.

Здесь минаретов купола - Дневные звёзды в небе ясном - Так чистая лазурь светла И вязь орнаментов прекрасна.

The domes of minarets stand proud - like day-time stars in clear sky shining: so pure the blue, without a cloud and fine the spires in art combining.

…  …  …

2/2

О, Самарканд! Ты - мудрость лет, Здесь рядом церкви и мечети… В тебя влюблённым лучше нет И краше города на свете!

Oh, Samarkand! You’re wise through age,

with mosques and churches closely neighboured.

For those who love you, the world’s stage has no town better nor more favoured.

О, Самарканд! Живи, цвети И славься добрыми делами! Навстречу вечности лети И вечно оставайся с нами.

Oh, Samarkand! Live long and thrive, be glorified by each good action! And while eternity’s your prize, be here our lasting benefaction.

Verkinto de tiu ĉi Rusa poemo estas Ирина Саверьевна Алексеева .

Arg-860-1733 (2014-05-15 12:22:00)

Tiu ĉi poemo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.-сост.: А. Ионесов . – Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил.

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en itala de Carlo Minnaja (∗1940-03-19).

Arg-860-2206 (2014-11-13 12:31:20)

Tiu ĉi poem-italigo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.-сост.: А. Ионесов . – Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил. Pri la tradukinto vidu la retejon http: // it. wikipedia. org/ wiki/ Carlo_Minnaja .

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en la Anglan de Malcolm Jones.

Arg-860-2037 (2014-10-16 20:52:12)

Tiu ĉi poem-angligo aperis en 2013 en libro “Oда Самарканду” (= Odo por Samarkando). Sinjoro Malcolm Jones, kiu angligis la poemon, estas brita esperantisto. Lia adreso troviĝas en la UEA-jarlibro de la jaro 2014 sur paĝo 126.

http://www.poezio.net/version?poem-id=860&version-id=1733,2206,2037